If a new machine can make 10,000 aluminum cans three times faster than an older machine. With both machines working 10,000 cans can be made in 9 hours. How long would it take the new machine, working alone, to make the 10,000 cans.
Answer by richwmiller(17219)

1/a+1/b=1/x
a=b*3
1/b*3+1/b=1/9
1/b*3+3/3*b=1/9
4/b*3=1/9
3*b=36
b=12
a=12*3
a=36 b=12
check
1/a+1/b=1/x
1/36+1/12=1/9
0.27777778e-1+0.83333333e-1=1/9
0.11111111=1/9
0.11111111=0.11111111
x=9 hours working together
a=36 hours working alone
b=12 hours working alone
ok
codewtr